Home‎ > ‎Crew 133‎ > ‎

108-CommandersReport

Commander Report
01/08/2014
Paula Crock
 
When I was reading through their backgrounds as we were pooling information for our application I
saw a crew able to handle the challenges that as a veteran I knew awaited us at MDRS should we be
selected.  Today my fellow UND students and our late-addition M.D. from Lille 2 University truly
proved themselves.  Jessup lead an EVA team with Komorowski, Howell and Gartrelle on what was
expected to be a short, easy task of setting up the radio telescope just outside the Hab followed by
collection of data at three locations for a GIS study.  This was only their second EVA so they
hadn’t yet had much time to acclimate to working in EVA gear nor much time to orient to the area.
The radio telescope setup proved to be more difficult than expected, complicated by frozen ground,
thick gloves, fogging helmets, make-shift tools and small parts but I could hear them over the radio
working together to solve their problems effectively and with good humor.  It took a tiring two
hours but they succeeded in getting the telescope set up.
 
That wasn’t enough for this crew, though!  After a brief break back in the Hab for water and snacks
they returned through the airlock to navigate through rugged Martian terrain north, east and south
of the Hab to collect data for the GIS study Jessup is working on while at MDRS.  Again through
radio chatter I could hear them working together to find better routes through the muddy areas and
rougher terrain.  All total, a strenuous 4 hour 40 minute EVA with two successful projects completed
and, as Gartrelle stated on his return, a better appreciation for how grueling the 7 or 8 hour EVAs
on ISS are.  Diaz-Rubin and I, who remained at the Hab today, have been promised a terrific photo
show of their EVA work after comms tonight, which is some consolation for having missed out on
getting to help out.


--------------------

CommanderCheckIn Report
01/08/2014
Paula Crock
Crew Physical Health:
Crewmember with sore back is able to perform duties but is still sore.
The rest of the crew are well.
  
Eva Depart Time:
09:43
Eva Return Time:
14:01
Narrative Of Field Mission Results:
Erected radio antenna; collected data on 3 points for Jessup project
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Hab Maint:
Power was out from 14:41 - 15:53.  Call was put in to D.G. who restarted
backup generator and noted that primary generator had faulty oil pressure
sensor.
Rep Sched:
Commander Check-In
Engineering Report
HSO Report
Science Report (GreenHab)
EVA Report
Daily Photos
Journalist Report
Commander Report
Tomorrow Plan:
We will all be in the Hab working on projects.
Inventory:
Still low on TP and trash bags
Support Req:
N/A
Miscelleneous:
N/A